Bob is twice as old as chuck,and susie is 10 years younger than bob.If the same of their ages is 100 years,how do is susie?
Answer by addingup(3677)

C = 1/2B
S = B-10
and
B+C+S = 100 Substitute:
B+1/2B+B-10 = 100
B+1/2B+B = 110
2.5B = 110
B = 44 Bob is 44
Susie is 44-10 = 34
Chuck is 1/2(44) = 22
:
John